OpenStudy (anonymous):

~PLEASE HELP ANYONE????Suppose you have a $15 coupon that can be used on any purchase over $100 at your favorite clothing store. The discounted price of a sports jacket is $203. What is the retail price of the jacket? a. $235.70 b. $303.00 c. $188.00 d. $218.00

OpenStudy (anonymous):

is $203 the price after the coupon? @jdoe0001

OpenStudy (anonymous):

@joemath314159

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

$203+($203*.15)

OpenStudy (anonymous):

i think let me dbl check

OpenStudy (anonymous):

there isnt an answer choice with 203,

OpenStudy (anonymous):

OP-(OP*off)=DP op-(op*.15)=203

OpenStudy (anonymous):

oh crap i was doing it wrong i read that as 15% off not 15 $ :P sorry

OpenStudy (anonymous):

203+15

OpenStudy (anonymous):

218 D

OpenStudy (anonymous):

The depth (inches) of water that accumulates in the spring soil from melted snow can be determined by dividing the cumulative winter snowfall (inches) by 12. Determine the amount of winter snow that accumulates 6 inches of water in the soil. a. 6 inches of snow b. 84 inches of snow c. 72 inches of snow d. 12 inches of snow

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

AW=AS/12 6=AS/12 6*12=AS

OpenStudy (anonymous):

72

OpenStudy (anonymous):

6*12 is 72 indeed